# Bifurcations and canards in the FitzHugh-Nagumo system: a tutorial in fast-slow dynamics

### v1

@@ Abstract.
In this article, we study the FitzHugh-Nagumo $(1,1)$–fast-slow system where the vector fields associated to the slow/fast equations come from the reduction of the Hodgin-Huxley model for the nerve impulse. After deriving dynamical properties of the singular and regular cases, we perform a bifurcation analysis and we investigate how the parameters (of the affine slow equation) impact the dynamics of the system. The study of codimension one bifurcations and the numerical locus of canards concludes this case-study. All theoretical results are numerically illustrated.
@@ 1. Introduction
Many animals have cells in their various physiological systems, such as the nervous, muscular, and cardiac systems, that are sensitive to certain electrical stimuli. These cells remain at rest most of the time, react to electrical stimulation at a given moment, and then return to a resting state until they are stimulated again. This oscillatory and excitability dynamics is present in neuronal activity and cardiac rhythm – see for instance Keener & Sneyd [12].
@@
In 1960–1961 Richard FitzHugh [4, 5] created a simplified version of the Hodgkin-Huxley model, considering only two variables. His goal was to simplify the HH model in order to make the dynamics of excitability more perceptible. The system developed by FitzHugh in 1961 has the following form:
@@
where $a\in{\mathbb{R}}\backslash\{0\},b,c\in{\mathbb{R}}$ are constants and $I$ may be seen as the intensity of a stimulus applied to the axon membrane. Variables $x$ and $y$ match with the pairs of variables $\left(v,m\right)$ and $\left(h,n\right)$ from the HH model, respectively.
During the 1960s, Jin-Ichi Nagumo constructed an electrical circuit that recreated FitzHugh’s system [18]. Nowadays, due to the contributions of both researchers, this model is known as the FitzHugh-Nagumo (FH-N) model. This model has been extensively studied see, for instance, [1] and references therein.
@@
Our goal with the present article is to exhibit the various dynamics in the multiple time scale model and motivate their existence in the light of methods from geometric singular perturbation theory and bifurcation theory.
@@ 2. Multiple Time Scale Theory and the FitzHugh-Nagumo System
In this section, general concepts of two time scale systems are introduced and illustrated with the FitzHugh-Nagumo system. Further results on fast-slow systems may be found in [16]. We start by introducing some necessary definitions.
@@
The fast time scale is represented by $t$ while $\tau$ refers to the slow time scale. In order to better understand the concepts and the terminology associated with these systems, the following definitions and results are illustrated by the FitzHugh-Nagumo (1,1)–fast-slow system.
@@
As described in [2], consider the FitzHugh-Nagumo (1,1)–fast-slow system:
@@
Both cubic and affine functions of Example 2.1, $f$ and $g$, come naturally from the reduction of the Hodgkin-Huxley model [12]. Note that in FH-N the functions $f$ and $g$ do not depend on $\varepsilon$. System (LABEL:fhn) will be referred to as the FitzHugh-Nagumo system in two variables.
3. The Singular Case ${\left(\varepsilon=0\right)}$
When analysing a fast-slow system, it is useful to consider the singular case $\varepsilon=0$.
@@ 3. The Singular Case ${\left(\varepsilon=0\right)}$
The critical manifold for the FitzHugh-Nagumo system (LABEL:fhn) is
@@ 2.       $\bullet$
$S_{0}=\mathcal{C}_{0\text{L}}\cup\mathcal{C}_{0\text{M}}\cup\mathcal{C}_{0\text{R}},$
@@
As expected, this explicit form of the slow subsystem is not defined at the fold points $x=\pm\dfrac{2}{\sqrt{3}}$.
@@
For $b>0$, if $|x^{*}|>\dfrac{2}{\sqrt{3}}$, the equilibrium attracts the slow flow. Otherwise, if $|x^{*}|<\dfrac{2}{\sqrt{3}}$ the equilibrium may either attract or repel the slow flow depending on the value of $1-b\left(4-3{x^{*}}^{2}\right)$ being negative or positive, respectively. Thus, the equilibrium point $\left(x^{*},y^{*}\right)$ of (LABEL:fhn) is stable if $\left(x^{*},y^{*}\right)\in\mathcal{C}_{0\text{L}}\cup\mathcal{C}_{0\text{R}}$ and unstable if $\left(x^{*},y^{*}\right)\in\mathcal{C}_{0\text{M}}$.
@@
In Example 3.2, we have disregarded the case where a fold point is also an equilibrium of the system, i.e., $x^{*}=\pm\dfrac{2}{\sqrt{3}}$. This particular situation requires special attention and will be analysed later in Section 6. For the remainder of this section, we describe the dynamics in the cases where all equilibria lie in $S_{0}$.
@@ 3.1. Dynamics
Figure 2. Examples of Cases (i)–(iv). Critical manifold in red - attracting regions in solid line and repelling region in dashed line. Line $\dot{y}=0$ in blue. Fast trajectories in yellow and slow trajectories in green. Points of interest/equilibrium/fold in black. Initial point, $A$, outlined in black.
Scenario (i) is trivial. If $B=\left(x_{b},y_{a}\right)$ is an equilibrium of (LABEL:fhn), then by definition $\left(\dot{x},\dot{y}\right)=0$, and therefore, the trajectory initiated at $A$ accumulates at $B$.
Scenario (ii) is not much different. Upon reaching the point $B$, the trajectory $\gamma_{A}$ moves through the slow flow of the system defined by $\dot{y}=x-by-c$. Since the equilibrium point $C$ is attracting, and since both points $B$ and $C$ are in $\mathcal{C}_{0\text{L}}$, the trajectory $\gamma_{A}$ moves from $B$ to $C$ along the branch $\mathcal{C}_{0\text{L}}$ of the critical manifold, where it accumulates.
@@
Scenario (iv): let $E=\left(x_{e},y_{e}\right)$ be an equilibrium point in $\mathcal{C}_{0\text{R}}$. Geometrically, the point $E$ is the intersection of the cubic curve $x^{\prime}=0$ with the line $\dot{y}=0$. Since there are no more equilibria in $\mathcal{C}_{0\text{L}}$, we may observe that, if $E$ is unstable, then $y_{a}>y_{e}$. The proof is simple: if $E$ is unstable, then $\psi_{x}<0$ where $\psi$ was defined in (10). Since $f_{x}\left(E,0\right)>0$, it follows that $b<0$ and therefore the line has a negative slope. Now, assume by contradiction that $y_{b}<y_{e}$. The line $\dot{y}=0$ necessarily intersects the cubic surface in $\mathcal{C}_{0\text{L}}$ and hence, there exists an equilibrium in $\mathcal{C}_{0\text{L}}$ giving our contradiction. Thus, at $B$, we have $\dot{y}>0$ and therefore, the trajectory $\gamma_{A}$ remains in $\mathcal{C}_{0\text{L}}$, with the first coordinate of $\gamma_{A}\left(\tau\right)$ going to $\infty$- ∞ as $\tau\rightarrow\infty$ while the second coordinate of $\gamma_{A}\left(\tau\right)$ goes to $\infty$. If the point $E$ is stable, $\gamma_{A}$ travels along $\mathcal{C}_{0\text{L}}$ from $B$ to $P_{-}$, jumps in fast time to point $D$ in the branch $\mathcal{C}_{0\text{R}}$, ending at $E$.
@@ 3.2. Period of the limit cycle
In this section we have described the dynamics of (LABEL:fhn) in the singular case $\varepsilon=0$, in the various cases according to the number and stability of its equilibria, as shown in Figures 2 and 3. We have also obtained conditions under which (LABEL:fhn) with $\varepsilon=0$ has a periodic solution that alternates between the fast and the slow flows and we have determined its period. In the next section we look at the dynamics of (3) in the nonsingular case $\varepsilon\overset{>}{\sim}0$.
4. The Regular Case ${\left(0<\varepsilon\ll 1\right)}$
@@
As we will see in Fenichel’s theorem (Theorem 1 stated below) when $0<\varepsilon\ll 1$, the dynamics of the perturbed system is similar to that of the singular system, with a deviation of $\mathcal{O}\left(\varepsilon\right)$, where $\mathcal{O}$ stands for the usual Landau notation. That is, the smaller the $\varepsilon$, the more similar the system trajectories are to those described in the singular system. Fenichel’s theorem guarantees that for $0<\varepsilon\ll 1$, there exists a set $S_{\varepsilon}$ very close to $S_{0}$ that exhibits the same behaviour as $S_{0}$.
@@ 4.1. The manifold $S_{\varepsilon}$
In what follows, our aim is to show how to compute the term $h_{1}(y)$ in the expansion. Taking into account that $g\left(x,y,\varepsilon\right)\neq 0$, write (14) as
@@
Since we defined the slow manifold $S_{\varepsilon}$ as an asymptotic expansion starting at the critical manifold $\mathcal{C}_{0}$, we can apply the same reasoning to determine the period of a periodic solution, $\mathcal{T}_{\gamma_{\varepsilon}}$, of a fast-slow system, with $0<\varepsilon\ll 1$. In particular, using the same argument used to prove (12) we have:
@@ 5. Bifurcations
with $0<\varepsilon\ll 1$. It has only one equilibrium $\mathrm{E}=\left(c,\varphi(c)\right)$ for $\varphi(x)\coloneqq 4x-x^{3}$, which results from the intersection of the cubic curve $y=\varphi(x)$ and the line $x=c$.
@@
When $c=\mathrm{c^{H}}\coloneqq\pm\dfrac{2}{\sqrt{3}}$ and $0<\varepsilon\ll 1$, we get $\lambda_{\mathrm{E}}^{\pm}\left(\mathrm{c^{H}},\varepsilon\right)=\pm\mathrm{i}\sqrt{\varepsilon}$. For $\varepsilon>0$, as shown in Figure 5, the stability of the equilibrium point in the neighbourhood of $\mathrm{c^{H}}$ changes from a stable focus to an unstable focus as $c$ decreases.
For $c=1.15<\mathrm{c^{H}}$ and $\varepsilon=1$, the equilibrium point $\mathrm{E}$ is an unstable focus, and thus it is expected that trajectories starting close to the fold point $P_{+}$ diverge, for $t>0$. Numerically, we found a periodic solution when $c\overset{<}{\approx}\mathrm{c^{H}}$ (cf. Figure 6(c)). This dynamics is typical of a subcritical Hopf bifurcation, with a stable periodic solution as described in the next result. Here subcritical means that the periodic solution exists for $c<c^{\mathrm{H}}$. If the periodic solutions were found for $c>c^{\mathrm{H}}$, the bifurcation would be called supercritical.
@@
Figure 5. Variation of the real and imaginary parts of the eigenvalues of the Jacobian matrix of system (22), evaluated at the equilibrium point $E$, as a function of the parameter $c$ ($\lambda_{\varepsilon}^{+}$ with solid line and $\lambda_{\varepsilon}^{-}$ with dashed line) for different values of $\varepsilon$, $\varepsilon=1$ (yellow), $\varepsilon=0.1$ (violet) e $\varepsilon=0$ (blue). Hopf bifurcation (black).
Figure 6. (a) - Bifurcation diagram of the parameter $c$ in the system (22), with $\varepsilon=0.5$. Stable equilibrium (solid line at $\mathcal{A}=0$) for $|c|>2/\sqrt{3}$, unstable (dashed line) for $|c|<2/\sqrt{3}$ and Hopf bifurcation points (black dots) for $|c|=2/\sqrt{3}$. Stable periodic solutions (solid line at $\mathcal{A}\neq 0$) for $|c|<2/\sqrt{3}$. (b) - Zoom of the region within the red rectangle observed in (a). (c) - Example of a periodic solution of the system (22) for $c=1.152$ and $\varepsilon=0.5$. $\mathcal{C}_{0}$ (red), vertical lines $x=c$ (blue) and $x=2/\sqrt{3}$ (black). Approaching/leaving trajectories from the equilibrium point (yellow/green) and periodic solution (violet). (d) - Zoom of the region within the magenta rectangle observed in (b).
Theorem 3 (Hopf, 1942. See also Marsden & McCraken, 1976 [17] for English translation in Section 5.).
@@
System (22) satisfies the conditions of Theorem 3, thus confirming the numerical finding of Figure 6(c). The stability of the periodic solution will be discussed in Section 6 below.
@@
with $0<\varepsilon\ll 1$. For all $b\in{\mathbb{R}}\setminus{\left\{0\right\}}$, the origin, $\mathrm{E}_{0}=\left(0,0\right)$, is an equilibrium point the system (23). If either $b<0$ or $b>1/4$, the system has two symmetric equilibria at $x_{\pm}=\pm\sqrt{4-b^{-1}}$. They are:
@@
For $b<0$, the equilibrium $E_{0}$ is an unstable node/focus and $E_{-}$ and $E_{+}$ are saddles. For $40<b\leq 1/$ 4, there is a stable limit cycle and only the origin is an equilibrium point of the system, but it still exhibits the same unstable node/focus dynamics.
The equilibrium $\mathrm{E}_{0}$ becomes a saddle when $b>1/4$. Moreover, at $b=1/4$ two new equilibria $\mathrm{E}_{-}$ and $\mathrm{E}_{+}$ are created (cf. Figure 8(a)). Given the symmetry of $E_{\pm}$ in this case, the stability of both equilibrium points is the same. Thus, we will only study the stability of $\mathrm{E}_{+}$ by computing
@@
and therefore, by Theorem 3, there is a Hopf bifurcation. In the limit, as $\varepsilon\to 0$, we have $\mathrm{b}_{\varepsilon}^{\mathrm{H}}\to 3/8$ (cf. Figure 7). Moreover, for $b=3/8$, the fold points are equilibria of (23).
The numerical simulations shown in Figure 8 indicate that the Hopf bifurcation at $\mathrm{E}_{+}$ is supercritical, hence generating an unstable periodic solution. This means that all trajectories starting in the regions bounded by the periodic solutions are attracted to the equilibrium point $\mathrm{E}_{-}$ or $\mathrm{E}_{+}$, also lying in the interior of that region. On the other hand, when trajectories start outside these regions, they converge to the global stable limit cycle which already existed when the equilibria were unstable. As observed in the example of Figure 8, the amplitude of the unstable periodic solutions grows very quickly for small variations in $b$, but reaches an unexpected maximum for $b=b_{\varepsilon}^{Hom}$. This happens when the periodic solution captures the saddle equilibrium $\mathrm{E}_{0}$ at a homoclinic bifurcation [8, 19].
@@
We already knew that (23) has the saddle equilibrium $\mathrm{E}_{0}=(0,0)$ and we have observed numerically that, for $b=b_{\varepsilon}^{Hom}$, there is a homoclinic trajectory to the origin $\mathrm{E}_{0}$. In Figures 8 and 9, which consider the system (23) for $\varepsilon=0.5$, the homoclinic orbit appears for $b=b_{\varepsilon}^{Hom}\approx 0.36932$, very close to $b_{\varepsilon}^{H}\approx 0.3666$. This bifurcation where a periodic solution is destroyed at a homoclinic orbit is called a homoclinic bifurcation [8, 19].

### v3

@@ Abstract.
In this article, we study the FitzHugh-Nagumo $(1,1)$–fast-slow system where the vector fields associated to the fast/slow equations come from the reduction of the Hodgin-Huxley model for the nerve impulse. After deriving dynamical properties of the singular and regular cases, we perform a bifurcation analysis and we investigate how the parameters (of the affine slow equation) impact the dynamics of the system. The study of codimension one bifurcations and the numerical locus of canards concludes this case-study. All theoretical results are numerically illustrated.
@@ 1. Introduction
Many animals have cells in their various physiological systems, such as the nervous, muscular, and cardiac systems, that are sensitive to certain electrical stimuli. These cells remain at rest most of the time, react to electrical stimulation at a given moment, and then return to a resting state until they are stimulated again. This oscillatory and excitability dynamics is present in neuronal activity and cardiac rhythm – see for instance [20].
@@
In 1960–1961 Richard FitzHugh [8, 9] created a simplified version of the HH model, considering only two variables. His goal was to simplify the HH model in order to make the dynamics of excitability more perceptible. The system developed by by FitzHugh in 1961 has the following form:
@@
where $a\in{\mathbb{R}}\backslash\{0\},b,c\in{\mathbb{R}}$ are constants and $I$ may be seen as the intensity of a stimulus applied to the axon membrane. Variables $x$ and $y$ of (1) describe the fast evolution of the membrane voltage of a neuron as well as the activation of sodium channels, and the inactivation/activation of the sodium/potassium channels, respectively. They match with the pairs of variables $\left(v,m\right)$ and $\left(h,n\right)$ from the HH model, respectively – cf. [9, 20].
During the sixties, Jin-Ichi Nagumo constructed an electrical circuit that recreated FitzHugh’s system [28]. Nowadays, due to the contributions of both researchers, this model is known as the FitzHugh-Nagumo (FH-N) model. This model has been extensively studied see, for instance, [2] and references therein.
@@
1.1. Literature
In this section, we review the FH-N equation, as well as its derivations, as one paradigmatic example of mathematical modelling. Although it has been motivated by neuroscience, the FH-N model exists in many variations of the original system and has been studied in several fields due to its simplicity and rich dynamics. Motivated by its broad applicability and generic dynamics, the FH-N system has triggered a huge body of research works characterising their bifurcations and the associated dynamical regimes.
There is an abundance of references in the literature. In what follows, we choose to mention only a few for clarity and the choice is based on our personal preferences. The reader interested in further detail and/or more examples can use the references within those we mention specially those in the review [2].
At the first glance, the simplifications introduced in the passage from the sophisticated HH model to the FH-N model might have a price. Phillipson and Schuster [29] explore numeric and analytically the four-dimensional Hodgkin–Huxley model and demonstrate that, under precise conditions, the FH-N equations can provide quantitative predictions in close agreement with the classic HH equations.
The bifurcations of FH-N system have been extensively analysed as a fast-slow system in [3, 11, 24, 21, 25, 4, 7, 22] uncovering complex patterns of behaviours that emerge from the simple and/or coupled FH-N model. They reveal how changes in the parameters can lead to different dynamical regimes, including equilibria, limit cycles, canards and chaos. Stability techniques allow the identification of critical points where small perturbations may lead to qualitatively different behaviours, shedding light on the robustness and sensitivity of excitable systems. Systems exhibiting excitability often display a threshold behaviour, where a certain level of input of perturbation is required to trigger the response – cf. [6].
We would like to refer the paper [10] for the study, not using the fast-slow approach, of (1), where global bifurcations are presented in the context of the global bifurcation diagram. Codimension one and two bifurcations of Hopf, homoclinic, saddle-node and Bogdanov-Takens are obtained.
The work [1] investigates the complex phenomena of canard explosion with mixed-mode oscillations, which can be observed in a fractional-order FH-N model. The study has highlighted the appearance of patterns of solutions with an increasing number of small-amplitude oscillations in each of such patterns, as one parameter of the fractional-order FH-N model is varied.
The paper [26] extends the FH-N model by including a time-varying threshold between electrical “silence” and electrical firing, as well as a periodic forcing for the intensity of the stimulus. Authors provide sufficient conditions for the existence of periodic solutions in such differential systems. This may be related with the work [32], where the author explores the dynamical behaviour of non-autonomous, almost-periodic discrete FH-N system defined on infinite lattices. The non-autonomous infinite-dimensional system has a uniform attractor which attracts all solutions uniformly with respect to the translations of external terms. Authors establish the upper semi-continuity of uniform attractors when the infinite-dimensional system is approached by a family of finite-dimensional systems.
In [5] authors propose a new method for the identification of the FH-N model dynamics via deterministic learning (with diffusion). The FH-N model is then described by partial differential equations which are transformed into a set of ordinary differential equations whose dynamics has been studied. Authors found spiral waves and an accurate identification of dynamics underlying the recurrent trajectory corresponding to any spatial point is achieved.
Spiking and bursting patterns observed in nerve membranes seem to be important when we investigate information representation model in the brain. Many topologically different bursting responses are observed in the mathematical models and their related bifurcation mechanisms have been clarified in [31]. In this article, the authors propose a design method to generate bursting responses in FitzHugh–Nagumo model with a simple periodic external force based on bifurcation analysis. Some effective parameter perturbations for the amplitude of the external input are given from the two-parameter bifurcation diagram.
1.2. Novelty and structure of the article
Our goal with the present article is to exhibit the various dynamics in the multiple time scale model and motivate their existence in the light of methods from geometric singular perturbation theory and bifurcation theory. This article studies the dynamics of a fast-slow system derived from the FH-N model according to the geometric singular perturbation theory and the bifurcation theory methods. We provide an analytic proof that the fast-slow FH-N system presents relaxation oscillation dynamics as well as periodic solutions induced by Hopf bifurcation – we fully analyse the type of bifurcations and we check rigorously all conditions for the emergence of a homoclinic orbit and canards connecting these two distinct dynamics. We illustrate each result with numerical computations.
@@ 2. Multiple Time Scale Theory and the FitzHugh-Nagumo System
In this section, general concepts of two time scale systems are introduced and illustrated with the FH-N system. Further results on fast-slow systems may be found in [25]. We start by introducing some necessary definitions.
@@
The fast time scale is represented by $t$ while $\tau$ refers to the slow time scale. In order to better understand the concepts and the terminology associated with these systems, the following definitions and results are illustrated by the FH-N (1,1)–fast-slow system.
@@
As described in [6], consider the FH-N (1,1)–fast-slow system:
@@
Both cubic and affine functions of Example 2.1, $f$ and $g$, come naturally from the reduction of the HH model [20]. Note that in FH-N the functions $f$ and $g$ do not depend on $\varepsilon$. System (LABEL:fhn) will be referred to as the FH-N system in two variables.
3. The Singular Case
When analysing a fast-slow system, it is useful to consider the singular case $\varepsilon=0$. Techniques such as nullclines analysis provide a geometric understanding of the model’s behaviour.
@@ 3. The Singular Case
The critical manifold for the FH-N system (LABEL:fhn) is
@@
(10) $S_{0}=\mathcal{C}_{0\text{L}}\cup\mathcal{C}_{0\text{M}}\cup\mathcal{C}_{0\text{R}},$
@@
As expected, this explicit form of the slow subsystem is not defined at the fold points $x=\pm 2/\sqrt{3}$.
@@
For $b>0$, if $|x^{*}|>2/\sqrt{3}$, the equilibrium attracts the slow flow. Otherwise, if $|x^{*}|<2/\sqrt{3}$ the equilibrium may either attract or repel the slow flow depending on the value of $1-b\left(4-3{x^{*}}^{2}\right)$ being negative or positive, respectively. Thus, the equilibrium point $\left(x^{*},y^{*}\right)$ of (LABEL:fhn) is stable if $\left(x^{*},y^{*}\right)\in\mathcal{C}_{0\text{L}}\cup\mathcal{C}_{0\text{R}}$ and unstable if $\left(x^{*},y^{*}\right)\in\mathcal{C}_{0\text{M}}$.
@@
In Example 3.2, we have disregarded the case where a fold point is also an equilibrium of the system, i.e., $x^{*}=\pm 2/\sqrt{3}$. This particular situation requires special attention and will be analysed later in Section 6. For the remainder of this section, we describe the dynamics in the cases where all equilibria lie in the set $S_{0}$ defined in (10).
@@ 3.1. Dynamics
Figure 2. Examples of Cases (i)–(iv). Critical manifold in red, attracting regions in solid line and repelling region in dashed line. Line $\dot{y}=0$ in blue. Fast trajectories in yellow and slow trajectories in green. Points of interest/equilibrium/fold in black. Initial point, $A$, outlined in black.
Scenario (i) is trivial. If $B=\left(x_{b},y_{a}\right)$ is an equilibrium of (LABEL:fhn), then by definition $\left(\dot{x},\dot{y}\right)=0$, and therefore, the trajectory initiated at $A$ accumulates at $B$ (cf. Figure 2(a)).
Scenario (ii) is not much different. Upon reaching the point $B$, the trajectory $\gamma_{A}$ moves through the slow flow of the system defined by $\dot{y}=x-by-c$. Since the equilibrium point $C$ is attracting, and since both points $B$ and $C$ are in $\mathcal{C}_{0\text{L}}$, the trajectory $\gamma_{A}$ moves from $B$ to $C$ along the branch $\mathcal{C}_{0\text{L}}$ of the critical manifold, where it accumulates (cf. Figure 2(b)).
@@
Scenario (iv): let $E=\left(x_{e},y_{e}\right)$ be an equilibrium point in $\mathcal{C}_{0\text{R}}$. Geometrically, the point $E$ is the intersection of the cubic curve $x^{\prime}=0$ with the line $\dot{y}=0$. Since there are no more equilibria in $\mathcal{C}_{0\text{L}}$, we may observe that, if $E$ is unstable, then $y_{a}>y_{e}$. The proof is simple: if $E$ is unstable, then $\psi_{x}<0$ where $\psi$ was defined in (11). Since $f_{x}\left(E,0\right)>0$, it follows that $b<0$ and therefore the line has a negative slope. Now, assume by contradiction that $y_{b}<y_{e}$. The line $\dot{y}=0$ necessarily intersects the cubic surface in $\mathcal{C}_{0\text{L}}$ and hence, there exists an equilibrium in $\mathcal{C}_{0\text{L}}$ giving our contradiction. Thus, at $B$, we have $\dot{y}>0$ and therefore, the trajectory $\gamma_{A}$ remains in $\mathcal{C}_{0\text{L}}$, with the first coordinate of $\gamma_{A}\left(\tau\right)$ going to $\infty$- ∞ as $\tau\rightarrow\infty$ while the second coordinate of $\gamma_{A}\left(\tau\right)$ goes to $\infty$ as in Figure 2(f). If the point $E$ is stable, $\gamma_{A}$ travels along $\mathcal{C}_{0\text{L}}$ from $B$ to $P_{-}$, jumps in fast time to point $D$ in the branch $\mathcal{C}_{0\text{R}}$, ending at $E$ (cf. Figure 2(e)).
@@ 3.2. Period of the limit cycle
In this section we have described the dynamics of (LABEL:fhn) in the singular case $\varepsilon=0$, in the various cases according to the number and stability of its equilibria, as shown in Figures 2 and 3. We have also obtained conditions under which (LABEL:fhn) with $\varepsilon=0$ has a periodic solution that alternates between the fast and the slow flows and we have determined its period. In the next section we look at the dynamics of (3) in the nonsingular case, i.e., when $\varepsilon\overset{>}{\sim}0$.
4. The Regular Case
@@
As we will see in Fenichel’s theorem (Theorem 1 stated below) when $0<\varepsilon\ll 1$, the dynamics of the perturbed system is similar to that of the singular system, with a deviation of $\mathcal{O}\left(\varepsilon\right)$, where $\mathcal{O}$ stands for the usual Landau notation. That is, the smaller the $\varepsilon$, the more similar the system trajectories are to those described in the singular system (cf. Figure 4(a)). Fenichel’s theorem guarantees that for $0<\varepsilon\ll 1$, there exists a set $S_{\varepsilon}$ very close to $S_{0}$ that exhibits the same behaviour as $S_{0}$.
@@
In what follows, our aim is to show how to compute the term $h_{1}(y)$ in the expansion. Taking into account that $g\left(x,y,\varepsilon\right)\neq 0$, we write (15) as
@@
Since we defined the slow manifold $S_{\varepsilon}$ as an asymptotic expansion starting at the critical manifold $\mathcal{C}_{0}$, we can apply the same reasoning to determine the period of a periodic solution, $\mathcal{T}_{\gamma_{\varepsilon}}$, of a fast-slow system, with $0<\varepsilon\ll 1$ (cf. Figure 4(b)). In particular, using the same argument used to prove (13) we have:
@@ 5. Bifurcations
with $0<\varepsilon\ll 1$. It has only one equilibrium $\mathrm{E}=\left(c,\varphi(c)\right)$ for $\varphi(x)\coloneqq 4x-x^{3}$, which results from the intersection of the cubic curve $y=\varphi(x)$ and the vertical line $x=c$.
@@
When $c=\pm\mathrm{c^{H}}\text{, with }\mathrm{c^{H}}=2/\sqrt{3}$, and $0<\varepsilon\ll 1$, we get $\lambda_{\mathrm{E}}^{\pm}\left(\mathrm{c},\varepsilon\right)=\pm\mathrm{i}\sqrt{\varepsilon}$. For $\varepsilon>0$, as shown in Figure 5, the stability of the equilibrium point in the neighbourhood of $\mathrm{c^{H}}$ changes from a stable focus to an unstable focus as $c$ decreases. Similarly, by symmetry, the dual stability transition occurs in the neighbourhood of $-\mathrm{c^{H}}$ as $c$ increases.
For $c=1.15<\mathrm{c^{H}}$ and $\varepsilon=1$, the equilibrium point $\mathrm{E}$ is an unstable focus, and thus it is expected that trajectories starting close to the fold point $P_{+}$ diverge, for $t>0$. Numerically, we found a periodic solution when $c\overset{<}{\approx}\mathrm{c^{H}}$ (cf. Figure 6(d)). This dynamics is typical of a subcritical Hopf bifurcation, with a stable periodic solution as described in the next result. Following the terminology of [12, Cap. IV, Section 2], subcritical means that the periodic solution exists for $c<c^{\mathrm{H}}$. If the periodic solutions are found for $c>c^{\mathrm{H}}$, the bifurcation would be called supercritical.
@@
Figure 5. Evolution of the real and imaginary parts of the eigenvalues of the Jacobian matrix of system (23), evaluated at the equilibrium point $E$, as a function of the parameter $c$ ($\lambda_{\varepsilon}^{+}$ with solid line and $\lambda_{\varepsilon}^{-}$ with dashed line) for different values of $\varepsilon$, $\varepsilon=1$ (green), $\varepsilon=0.5$ (yellow), $\varepsilon=0.1$ (blue) e $\varepsilon=0$ (violet). Hopf bifurcation (black).
Figure 6. (a), (b) and (c) Bifurcation diagrams: length ${\mathcal{A}}$ of the periodic solutions of system (23) with $\varepsilon=0.5$ as a function of the parameter $c$ with different levels of magnification. Equilibria correspond to ${\mathcal{A}}=0$. The graphs in (b) and (c) are zooms of the red and magenta rectangles in (a) and (b), respectively. Conventions: Stable periodic solutions (solid line at $\mathcal{A}\neq 0$) for $|c|<2/\sqrt{3}$. (d) Example of a periodic solution of the system (23) for $c=1.152$ and $\varepsilon=0.5$. $\mathcal{C}_{0}$ (red), vertical lines $x=c$ (blue) and $x=2/\sqrt{3}$ (black). Approaching/leaving trajectories from the equilibrium point (yellow/green) and periodic solution (violet).
Theorem 3 (Hopf, 1942 [19]. See also Marsden & McCraken, 1976 [27] for English translation in Section 5).
@@
System (23) satisfies the conditions of Theorem 3, thus confirming the numerical finding of Figure 6(d). The direction of bifurcation and the stability of the periodic solution will be discussed in more detail in Section 6 below.
@@
with $0<\varepsilon\ll 1$. For all $b\in{\mathbb{R}}\setminus{\left\{0\right\}}$, the origin, $\mathrm{E}_{0}=\left(0,0\right)$, is an equilibrium point the system (24). If either $b<0$ or $b>1/4$, the system has two symmetric equilibria at $x_{\pm}=\pm\sqrt{4-b^{-1}}$, as a result of a Pitchfork bifurcation, as the reader may check in Figure 8(a). They are:
@@
For $b<0$, the equilibrium $E_{0}$ is an unstable node/focus and $E_{-}$ and $E_{+}$ are saddles. For $40<b\leq 1/$ 4, there is a stable limit cycle (similar scenario to that of (v) represented in Figure 3(a)) and only the origin is an equilibrium point of the system, but it still exhibits the same unstable node/focus dynamics.
The equilibrium $\mathrm{E}_{0}$ becomes a saddle when $b>1/4$. Moreover, at $b=1/4$ two new equilibria $\mathrm{E}_{-}$ and $\mathrm{E}_{+}$ are created (cf. Figure 8(a)). Given that $E_{+}$ and $E_{-}$ are symmetric, the stability of both equilibrium points is the same. Thus, we will only study the stability of $\mathrm{E}_{+}$ by computing
@@
and therefore, by Theorem 3, there is a Hopf bifurcation. In the limit, as $\varepsilon\to 0$, we have $\mathrm{b}_{\varepsilon}^{\mathrm{H}}\to 3/8$ (cf. Figure 7).
Moreover, for $b=3/8$, the fold points are equilibria of (24).
The numerical simulations shown in Figure 8(c) indicate that the Hopf bifurcation at $\mathrm{E}_{+}$ is supercritical, hence generating an unstable periodic solution. This will be confirmed analytically in Section 6 below. This means that all trajectories starting in the regions bounded by the periodic solutions are attracted to the equilibrium point $\mathrm{E}_{-}$ or $\mathrm{E}_{+}$, also lying in the interior of that region. On the other hand, when trajectories start outside these regions, they converge to the global stable limit cycle which already existed when the equilibria were unstable. As observed in the example of Figure 8(c), the length (a numerical estimate of the perimeter of the curve) of the unstable periodic solutions grows very quickly for small variations in $b$, but reaches an unexpected maximum for $b=b_{\varepsilon}^{Hom}$. This happens when the periodic solution captures the saddle equilibrium $\mathrm{E}_{0}$ at a homoclinic bifurcation [15, 30].
@@
We already knew that (24) has the saddle equilibrium $\mathrm{E}_{0}=(0,0)$ and we have observed numerically that, for $b=b_{\varepsilon}^{Hom}$, there is a homoclinic trajectory to the origin $\mathrm{E}_{0}$. In Figures 8(c) and 9, which consider the system (24) for $\varepsilon=0.5$, the homoclinic orbit appears for $b=b_{\varepsilon}^{Hom}\approx 0.36932$, very close to $b_{\varepsilon}^{H}\approx 0.3666$. The bifurcation where a periodic solution is destroyed at a homoclinic orbit is called a homoclinic bifurcation [15, 30].